# devoted to memory of Alexei Alexandrovich Starobinsky

### v15

@@ II.2 Decay of composite particles in de Sitter spacetime
The same local temperature $T=H/\pi$ describes the process of the splitting of the composite particle with mass $m$ into two components with $m_{1}+m_{2}>m$, which is also not allowed in the Minkowski vacuum Bros2008 ; Bros2010 ; Jatkar2012 ; Volovik2009 . In the limit $m\gg H$, the rate of such decay of the composite particle is
$w\sim\exp{\left(-\frac{\pi(m_{1}+m_{2}-m)}{H}\right)}=\exp{\left(-\frac{m_{1}+m_{2}-m}{T}\right)}\,.$ (6)
In the same way, the particle with mass $m$ can radiate the pair of the other particles moving in the opposite directions with the rate
$w\sim\exp{\left(-\frac{\pi(3m-m)}{H}\right)}=\exp{\left(-\frac{2m}{T}\right)}\,.$ (7)
@@
For each of the two detectors, which detect the created particles moving in the opposite directions, the radiation rate looks as thermal with the Gibbons-Hawking temperature $T_{\rm GH}=T/2$. However, since two particles are created coherently, the temperature describing the radiation of the pair is twice the Gibbons-Hawking temperature. The same connection between the local temperature and the temperature of the Hawking radiation from the cosmological horizon is discussed in Sec. II.3.
@@
The similar processes take place in the so-called Cosmological Collider,Maldacena2015 ; Reece2023 where the new particle created by the Hawking radiation plays the role of the external object which produces the heavy particles. Here there are two different physical processes, which are described by different temperatures. The Hawking radiation from the de Sitter vacuum is determined by the Hawking temperature $T_{\rm GH}$ of the cosmological horizon, while the further process – the splitting of the created particles – is determined by the local temperature $T=2T_{\rm GH}$.
Note that the creation of new particles by the particle in the de Sitter environment demonstrates that the de Sitter vacuum is unstable towards the creation of matter. The consequence of such instability will be discussed in Sec. VII.
II.3 Connection between the local and Hawking temperatures
@@
The local temperature $T=H/\pi$ also determines the process of the Hawking radiation from the cosmological horizon and the Gibbons-Hawking temperature $T_{\rm GH}=T/2=H/2\pi$. The reason is that in the Hawking process, two particles are coherently created (the analog of cotunneling): one particle is created inside the horizon, while its partner is simultaneously created outside the horizon.Parikh2002 The rate of the coherent radiation of two particles, each with energy $E$, is $w\propto\exp(-\frac{2E}{T})$. However, the observer who uses the Unruh-DeWitt detector can detect only the particle created inside the horizon. For this observer the creation rate $w\propto\exp(-\frac{2E}{T})$ is perceived as
@@
On the contrary, in the local process of the decay of the atom, which is not related to the cosmological horizon, only single particle (electron) is radiated from the atom. This process is fully determined by the local temperature, $w\propto\exp(-\frac{\epsilon_{0}}{T})$.
II.4 Two detectors: exited atom vs ionized atom
Note the main difference between the temperature measured by the observer using the Unruh-DeWitt detector (see e.g. Ref.Conroy2022 and references therein) and the temperature measured by the observer using ionization of an atom. The ionization process is possible, because the radiated electron moves far away from the atom to position $r=r_{0}$, where its negative gravitational energy compensates the ionization potential. In the Unruh-DeWitt detector, which corresponds to the two-level atom interacting with a quantum field, the electron in the atom is excited but remains in the same position in the same atom. That is why such excitation of electron may only come either by the Hawking radiation, or by radiation of photons, see Sec. II.5.
However, in the de Sitter case there is no difference in temperatures measured by two detectors. If in the Unruh-DeWitt detector experiments the Hawking temperature is properly interpreted according to Eq.(8), the measured temperature of the Hawking radiation is also $T=2T_{\rm GH}$. So, in spite of different physical principles, both detectors in the de Sitter environment show the same physical temperature, $T=2T_{\rm GH}$. This also demonstrates the uniqueness of the de Sitter state.
II.5 Radiation of photons by atom in de Sitter environment
Instead of the radiation of electron from the atom, one can consider the radiation of photon by the same atom. In both cases the escape of photon or electron from the atom provides the negative gravitational energy, which allows to consider the processes which are prohibited in the Minkowski state. Let us consider the atom in the excited state with energy $E_{0}+\epsilon$, where $E_{0}$ is the ground state energy. In the de Sitter environment, the radiation of photon is possible even if its energy $cp$ exceeds the energy difference between the excited and the ground state level, i.e. when $cp>\epsilon$.
For the relativistic photon the equation (2) takes the form (here we use $c=1$):
$\sqrt{p_{r}^{2}+p_{\perp}^{2}}+p_{r}v(r)=\epsilon\,.$ (9)
This gives the trajectory of photon
@@ II.5 Radiation of photons by atom in de Sitter environment
Eq.(11) demonstrates, that radiation takes place even if the atom is in the ground state, i.e. when $\epsilon\rightarrow 0$. In this case Eq.(11) describes the process of radiation of photon in the de Sitter environment:
@@
In this process, the atom in its ground state serves as an external object, which violates the de Sitter symmetry and thus provides the nonzero matrix element for the radiation of photon.
II.6 Accelerating detector: is there connection between the local process and Unruh radiation?
In the same way as in Sections II.4 and II.5, one may consider different independent processes related to the accelerating detector. One of them is the Unruh radiation measured by the Unruh-DeWitt detector, which could be linked to the apparent event horizon in the Rindler spacetime and to the corresponding Unruh temperature $T_{U}=a/2\pi$, where $a$ is acceleration.Unruh1976 ; Unruh2019 ; Unruh2022 The other processes, such as the local process of ionization of the accelerating atom,Volovik2022e have no relation to the event horizon.
Let us consider the process of ionization of atom. Similar to Eq.(5), the trajectory $p_{x}(x)$ of the radiated electron is obtained from the classical equation
$\frac{{\bf p}^{2}}{2m}-max=-\epsilon_{0}\,.$ (13)
Here we used the equivalence between the acceleration of the reference frame and the constant gravitational field, $g=a$.
In the WKB approximation and for $a\ll\epsilon_{0}\ll m$ this gives the radiation rate:Volovik2022e
$w\sim\exp\left(-\frac{4\sqrt{2}}{3}\frac{\epsilon_{0}}{a}\left(\frac{\epsilon_{0}}{m}\right)^{1/2}\right)\,.$ (14)
In principle, one can introduce the effective temperature, which is similar to the effective temperature $T_{\rm eff}\sim El$ of hopping electrons in a strong electric field $E$, where $l$ is the localization length.Shklovskii1992 ; Shklovskii2024 In our case, the analog of the localization length is $l=1/\sqrt{m\epsilon_{0}}$. But otherwise the local process of ionization of the accelerated atom is certainly non-thermal. Since $\epsilon_{0}\ll m$, the rate of ionization in Eq.(14) essentially exceeds the ionization rate in any thermal process, which involves acceleration, $w_{\rm thermal}\sim\exp(-\gamma\epsilon_{0}/a)$. This is very different from the ionization of the atom in the de Sitter environment, which looks thermal.
In principle, with the same approach one can also calculate the Unruh radiation of photons using different arrangements of detectors. The state of the art in theoretical understanding of Unruh radiation can be found in Refs. Gregori2023 ; Gregori2022 ; Matsuda2024 and references therein. Our examples demonstrate that the rate of Unruh radiation depends on details of the considered processes and it is not necessarily thermal.
II.7 de Sitter symmetry and de Sitter heat bath
@@ II.6 Accelerating detector: is there connection between the local process and Unruh radiation?
The uniqueness of the de Sitter thermal state lies in the fact that its temperature does not violate the de Sitter symmetry, and thus does not require the preferred reference frame. This is distinct from thermal state of matter, which always has a preferred reference frame where matter is at rest. As a result, thermal matter violates the de Sitter symmetry, which leads to the heat exchange between the two thermal subsystems, gravity and matter.
II.8 From local temperature to local entropy
The gravity subsystem – de Sitter quantum vacuum – has its own temperature and entropy, see also Ref.Padmanabhan2020 . On the other hand the de Sitter spacetime serves as the thermal bath for matter. Then the quasi-equilibrium states of the expanding Universe can be described by two different temperatures: the temperature of the gravitational vacuum (temperature of dark energy) and the temperature of the matter degrees of freedom.Vergeles2023 In this section we discuss the pure de Sitter vacuum without the excited matter ignoring for the moment the thermal activated creation of matter from the vacuum. The excitation and thermalization of matter by the de Sitter thermal bath will be discussed in Sec. VII.
If the vacuum thermodynamics is determined by the local the activation temperature $T=H/\pi$, then in the Einstein gravity with cosmological constant the vacuum energy density is quadratic in temperature:
@@ II.8 From local temperature to local entropy
This leads to the free energy density of the de Sitter vacuum, $F=\epsilon_{\rm vac}-Td\epsilon_{\rm vac}/dT$, which is also quadratic in $T$, and thus the entropy density $s_{\rm vac}$ in the de Sitter vacuum is linear in $T$:
@@
The Fermi velocity $v_{F}$ and Fermi momentum $p_{F}$ of this cosmological analog of Fermi liquid are on the order of the speed of light and the inverse Planck length correspondingly, $v_{F}\sim c$ and $p_{F}\sim 1/l_{\rm Pl}$. Of course, this is only the parallel, and one should not identify the de Sitter vacuum with the real Fermi liquid. However, the real Fermi surface may form inside the black hole horizon.HuhtalaVolovik2002 ; Zubkov2019
II.9 Hubble volume entropy vs entropy of the cosmological horizon

### v16

@@ II.2 Decay of composite particles in de Sitter spacetime
The same local temperature $T=H/\pi$ describes the process of the splitting of the composite particle with mass $m$ into two components with the larger total mass, $m\rightarrow m_{1}+m_{2}>m$, which is also not allowed in the Minkowski vacuum.Bros2008 ; Bros2010 ; Jatkar2012 ; Volovik2009
It is instructive to derive the rate of this process using also the semiclassical tunneling picture. For simplicity we consider creation of particles with equal masses, $m_{1}=m_{2}>m/2$. The trajectory of each of the two particles with mass $m_{1}$ moving in the radial direction from the origin at $r=0$ is obtained from equation
$E(p_{r},r)=\sqrt{p_{r}^{2}+m_{1}^{2}}+p_{r}Hr=\frac{m}{2}~{}.$ (8)
We took into account that each of the two particles carries the one half of the energy of the original particle, i.e. $E=m/2$. The momentum along the trajectory is
$p_{r}(r)=\frac{1}{1-H^{2}r^{2}}\left[-\frac{m}{2}Hr+\sqrt{\frac{m^{2}}{4}-m_{1}^{2}+m_{1}^{2}H^{2}r^{2}}\right].$ (9)
This momentum is imaginary in the classically forbidden region $r<r_{0}$, where $r_{0}=\sqrt{1-m^{2}/4m_{1}^{2}}/H$. This gives the imaginary contribution to the action:
[…]
where as before $r_{H}=1/H$ is the position of the de Sitter horizon. We must take into account that due to momentum conservation, the two particles tunnel simultaneously in the opposite directions. Such a coherent process of co-tunneling adds a multiplier of 2 to the argument of the exponent. As a result one obtains the decay rate of the composite particle into two particles, which is again governed by the temperature $T=H/\pi$:
$\Gamma(1\rightarrow 2)\propto\exp(-4\,{\bf Im}S)=\exp\left(-\frac{\pi(2m_{1}-m)}{H}\right)~{}.$ (11)
In general case, when $m_{2}\neq m_{1}$, the rate of the decay of the composite particle in the limit $m\gg H$ is determined by the same temperature:
$\Gamma(1\rightarrow 2)\sim\exp{\left(-\frac{\pi(m_{1}+m_{2}-m)}{H}\right)}=\exp{\left(-\frac{m_{1}+m_{2}-m}{T}\right)}\,.$ (12)
II.3 Triplication of particles in de Sitter spacetime
In the same way, massive fermions can be reproduced in the de Sitter environment. For example, the electron at $r=0$ can create an electron-positron pair, with electron and positron moving in opposite directions. In this process the fermion is triplicated with the rate given by:
$\Gamma(1\rightarrow 3)\sim\exp{\left(-\frac{\pi(3m-m)}{H}\right)}=\exp{\left(-\frac{2m}{T}\right)}\,.$ (13)
Since the created electron and positron move in the opposite directions, they are detected by two different detectors. Each of the two detectors can detect only single particle, and thus for each detector the radiation rate looks as thermal with the Gibbons-Hawking temperature $T_{\rm GH}=T/2$:
$\Gamma(1\rightarrow 3)\sim\exp{\left(-\frac{2m}{T}\right)}=\exp{\left(-\frac{m}{T/2}\right)}=\exp{\left(-\frac{m}{T_{\rm GH}}\right)}\,.$ (14)
@@ II.3 Triplication of particles in de Sitter spacetime
However, in this process two particles are created simultaneously and coherently (the so-called co-tunneling process). That is why the temperature, which describes such co-tunneling process of radiation of the pair, is twice the Gibbons-Hawking temperature. The same connection between the local temperature and the temperature of the Hawking radiation from the cosmological horizon is discussed in Sec. II.4.
It is important that each of three particles (the original electron + electron-positron pair, which is created in the process in Eq.(13)) is able to create the other three particles, and then the triplication process continues further. After $n$ replications there will be $3^{n}$ fermions ($(3^{n}+1)/2$ electrons and $(3^{n}-1)/2$ positrons. Such creation of multiple particles from a single particle immersed in the de Sitter environment demonstrates that the de Sitter vacuum is unstable towards the creation of matter. The consequence of such instability will be discussed in Sec. VI.
@@
The similar processes take place in the so-called Cosmological Collider,Maldacena2015 ; Reece2023 where the new particle created by the Hawking radiation plays the role of the external object which produces the heavy particles. In this case we have two different physical processes: the Hawking radiation from the cosmological horizon, and the further local process – the splitting of the created particles, which is determined by the local temperature.
II.4 Connection between the local and Hawking temperatures
@@
The local temperature $T=H/\pi$ also determines the process of the Hawking radiation from the cosmological horizon and the Gibbons-Hawking temperature $T_{\rm GH}=T/2=H/2\pi$. The reason is that in the Hawking process, two particles are created coherently (this is the analog of cotunneling): one particle is created inside the horizon, while its partner is simultaneously created outside the horizon.Parikh2002 The rate of the coherent cotunneling of two particles, each with energy $E$, is $w\propto\exp(-\frac{2E}{T})$. However, the observer who uses the Unruh-DeWitt detector can detect only the particle created inside the horizon. For this observer with limited information, the creation rate $w\propto\exp(-\frac{2E}{T})$ is perceived as
@@
The similar scenario of doubling the Hawking temperature was suggested for the black hole horizon. Hooft2022 ; Hooft2023 In this case the partner of the created particle is in the mirror image of the space-time which replaces the coordinate singularity. Although the existence of such quantum clone of the black hole is problematic, the consideration of the coherent creation of two partners on the two sides of the cosmological horizon is applicable to the de Sitter vacuum.
On the contrary, in the local process of the decay of an atom, which is not related to the cosmological horizon, only single particle (electron) is radiated from the atom. This process is fully determined by the local temperature, $w\propto\exp(-\frac{E_{n}}{T})$.
II.5 Two detectors: exited atom vs ionized atom
Note the main difference between the temperature measured by the observer using the Unruh-DeWitt detector (see e.g. Ref.Conroy2022 and references therein) and the temperature measured by the observer using ionization of an atom. The ionization process is possible, because the radiated electron moves far away from the atom to position $r=r_{0}$, where its negative gravitational energy compensates the ionization potential. In the Unruh-DeWitt detector, which corresponds to the two-level atom interacting with a quantum field, the electron in the atom is excited but remains in the same position in the same atom. That is why such excitation of electron may only come either by the Hawking radiation, or by radiation of photons, see Sec. II.6.
However, in the de Sitter case there is no difference in temperatures measured by two detectors. If in the Unruh-DeWitt detector experiments, the observer properly interprets the result of the observations according to Eq.(15), the measured temperature of the Hawking radiation is also $T=2T_{\rm GH}$. So, in spite of different physical principles, both detectors in the de Sitter environment show the same physical temperature, $T=2T_{\rm GH}$. This also demonstrates the uniqueness of the de Sitter state with its symmetry under the generalized translations.
II.6 Radiation of photons by atom in de Sitter environment
Instead of the radiation of electron from the atom, one can consider the radiation of photon by the same atom. In both cases the escape of photon or electron from the atom provides the negative gravitational energy. That is why these processes, which are prohibited in the Minkowski spacetime, are energetically possible in the de Sitter environment. Let us consider the atom in the excited state with energy $E_{0}+\epsilon$, where $E_{0}$ is the ground state energy of the atom. In the de Sitter environment, the radiation of photon is possible even if its energy $cp$ exceeds the energy difference between the excited and the ground state level, i.e. when $cp>\epsilon$.
For the relativistic photon the energy conservation equation (3) takes the form (here we use $c=1$):
$\sqrt{p_{r}^{2}+p_{\perp}^{2}}+p_{r}Hr=\epsilon\,.$ (16)
This gives the following trajectory of photon:
@@ II.6 Radiation of photons by atom in de Sitter environment
Eq.(18) demonstrates, that radiation takes place even if the atom is in the ground state, i.e. when $\epsilon\rightarrow 0$. In this case Eq.(18) describes the process of radiation of a single photon in the de Sitter environment:
@@
In this process, the atom in its ground state serves as an external object, which violates the de Sitter symmetry and thus provides the nonzero matrix element for the radiation of photon. This process is certainly different from the Hawking process of radiation.
II.7 Accelerating detector: is there connection between the local process and Unruh radiation?
In the same way as in Sections II.5 and II.6, one may consider different independent processes related to the accelerating detector. One of them is the Unruh radiation measured by the Unruh-DeWitt detector, which could be linked to the apparent event horizon in the Rindler spacetime and to the corresponding Unruh temperature $T_{U}=a/2\pi$, where $a$ is acceleration.Unruh1976 ; Unruh2019 ; Unruh2022 The other processes, such as the local process of ionization of the accelerating atom,Volovik2022e have no relation to the Rindler horizon.
Let us consider the process of ionization of an atom moving with acceleration. Similar to Eq.(6), the trajectory $p_{x}(x)$ of the radiated electron is obtained from the classical equation
$\frac{{\bf p}^{2}}{2m}-max=-\epsilon\,.$ (20)
Here, we used the equivalence between the acceleration of the reference frame and the constant gravitational field, $g=a$; and $\epsilon$ is the ionization potential.
In the WKB approximation and in the limit of slow acceleration, $a\ll\epsilon\ll m$, the classically forbidden trajectory of electron escaping from the atom gives the radiation rate,Volovik2022e which is similar to the radiation rate in the electric field:LandauLifshitzQM
$w\sim\exp\left(-\frac{4\sqrt{2}}{3}\frac{\epsilon}{a}\left(\frac{\epsilon}{m}\right)^{1/2}\right)\,.$ (21)
In principle, one can introduce the effective temperature, which is similar to the effective temperature $T_{\rm eff}\sim El$ of hopping electrons in a strong electric field $E$, where $l$ is the localization length.Shklovskii1992 ; Shklovskii2024 In our case, the analog of the localization length is $l=1/\sqrt{m\epsilon}$. But otherwise the local process of ionization of the accelerated atom is certainly non-thermal. Since $\epsilon\ll m$, the rate of ionization in Eq.(21) essentially exceeds the ionization rate in any thermal process, which involves acceleration, $w_{\rm thermal}\sim\exp(-\gamma\epsilon/a)$. This is very different from the ionization of the atom in the de Sitter environment, which looks thermal.
With the same approach one can also calculate the Unruh radiation of photons using different arrangements of detectors. The state of the art in theoretical understanding of Unruh radiation can be found in Refs. Gregori2023 ; Gregori2022 ; Matsuda2024 and references therein. Our examples demonstrate that the rate of Unruh radiation depends on details of the considered processes and it is not necessarily thermal. Thermal behavior of the local and global processes in the de Sitter environment is the result of the special de Sitter symmetry.
III Thermodynamics of the de Sitter state
III.1 de Sitter symmetry and de Sitter heat bath
@@ III Thermodynamics of the de Sitter state
The uniqueness of the de Sitter thermal state lies in the fact that its temperature does not violate the de Sitter symmetry, and thus does not require the preferred reference frame. This is distinct from thermal state of matter, which always has a preferred reference frame where matter is at rest. As a result, thermal matter violates the de Sitter symmetry, which leads to the heat exchange between the two thermal subsystems, gravity and matter, and finally to the de Sitter decay.
III.2 From local temperature to local entropy
The gravity subsystem – de Sitter quantum vacuum – has its own temperature and entropy, see also Ref.Padmanabhan2020 . On the other hand the de Sitter spacetime serves as the thermal bath for matter. Then the quasi-equilibrium states of the expanding Universe can be described by two different temperatures: the temperature of the gravitational vacuum (temperature of dark energy) and the temperature of the matter degrees of freedom.Vergeles2023 In this section we discuss the pure de Sitter vacuum without the excited matter ignoring for the moment the thermally activated creation of matter from the vacuum. The excitation and thermalization of matter by the de Sitter thermal bath will be discussed in Sec. VI.
If the vacuum thermodynamics is determined by the local activation temperature $T=H/\pi$, then in the Einstein gravity with cosmological constant the vacuum energy density is quadratic in temperature:
@@ III.2 From local temperature to local entropy
This leads to the free energy density of the de Sitter vacuum, $F=\epsilon_{\rm vac}-Ts_{\rm vac}$, which is also quadratic in $T$, and thus the entropy density $s_{\rm vac}$ in the de Sitter vacuum is linear in $T$:
@@
III.3 de Sitter vacuum, Fermi liquid and cosmological constant problem
@@
The Fermi velocity $v_{F}$ and Fermi momentum $p_{F}$ of this cosmological analog of the Fermi liquid are on the order of the speed of light and the inverse Planck length correspondingly, $v_{F}\sim c$ and $p_{F}\sim M_{\rm Pl}=1/l_{\rm Pl}$. Of course, this is only a parallel, and one should not identify the de Sitter vacuum with a real Fermi liquid. Although the real Fermi surface may form inside the black hole horizon.HuhtalaVolovik2002 ; Zubkov2019
Since the thermodynamics of the de Sitter state with the thermal energy $\epsilon_{\rm vac}\propto T^{2}$ is similar to the thermodynamics of the Fermi liquid, let us try to exploit this connection. One of the directions is the Sommerfeld law in Fermi liquid, which states that the entropy per one atom of the Fermi liquid is
$S=\frac{s_{\rm FL}}{n_{\rm FL}}\sim\frac{T}{E_{F}}\,,$ (25)
where $n_{\rm FL}\sim p_{F}^{3}$ is the density of atoms in the Fermi liquid and $E_{F}$ is Fermi energy.
We do not know what are the ”atoms of the vacuum”, but from Eq.(23) it follows that the entropy density of the vacuum $s_{\rm vac}\sim T/l_{\rm Pl}^{2}\sim(T/M_{\rm Pl})/l_{\rm Pl}^{3}$, where $l_{\rm Pl}$ is the Planck length and $M_{\rm Pl}$ is the Planck energy. This suggests that the density of the ”atoms of the vacuum” is $n_{\rm Pl}\sim 1/l_{\rm Pl}^{3}$ and entropy per ”atom of the vacuum” is:
$S=\frac{s_{\rm vac}}{n_{\rm Pl}}\sim s_{\rm vac}l_{\rm Pl}^{3}\sim\frac{T}{M_{\rm Pl}}\,.$ (26)
Eq.(26) is the full analog of the Sommerfeld law for Fermi liquid. This analogy also suggests that the corresponding density of states in the quantum vacuum (the analog of density of states at the Fermi level $N_{F}\sim mp_{F}$ in Fermi liquids) is $N_{\rm Pl}\sim M_{\rm Pl}^{2}$. For bosonic and fermionic degrees of freedom of this quantum vacuum the density of states is $N_{\rm Pl}=9/4\pi G$ and $N_{\rm Pl}=9/2\pi G$ correspondingly, which can be compared with the value $N_{\rm Pl}=3\pi/G$ suggested in Ref. Chu2023 . This huge density of states leads to a very large entropy of the de Sitter state even at very small temperature of the vacuum.
So, the quantum vacuum looks as some specific form of the relativistic Fermi liquid. However, some consequences are the same as in the non-relativistic Fermi liquid. In particular, in the full equilibrium at $T=0$ and in the absence of external pressure, $P=0$, the energy density of the non-relativistic Fermi liquid is exactly zero, $\epsilon(n)-\mu n=Ts-P=0$. This follows solely from thermodynamics, and is valid for any macroscopic system in its ground state. In all such systems, the contribution of the zero-point energies of collective modes (phonons, fermionic quasiparticles, magnions, etc.) is exactly compensated by the microscopic (atomic) degrees of freedom. This thermodynamic property does not depend on the microscopic physics, and is easily checked in condensed matter systems, where we know both the micro and macro physics. That is why it is not surprizing, that for the relativistic quantum vacuum, the vacuum energy density is zero if the following conditions are fulfiled: there is the full thermodynamic equilibrium; the temperature $T=0$; there is no matter (no quasiparticles); and there is no external pressure, $\epsilon_{\rm vac}(H=0,T=0)=-P_{\rm vac}=0$. This demonstrates, that thermodynamics, which is insensitive to the microscopic structure of the vacuum (ground state), solves the main cosmological constant problem. The other cosmological constant problems are discussed in Sec. VI.6.
III.4 Hubble volume entropy vs entropy of the cosmological horizon